OpenStudy (anonymous):

Find the x-intercepts for the parabola defined by the equation below. y = 3x^2 + 18x + 15 so i would do x+3=0 and x+15=0 and solve did i do that right

OpenStudy (anonymous):

No...factor out the 3 from the equation leaving \[3(x^{2}+6x+5)\])

OpenStudy (anonymous):

Now factor the trinomial...

OpenStudy (anonymous):

3x^2 + 18x+5

OpenStudy (anonymous):

no + 15

OpenStudy (anonymous):

Once factored out, you can leave the 3 alone for awhile. What multiplies to 5 and adds to 6

OpenStudy (anonymous):

nothing..

OpenStudy (anonymous):

36?

OpenStudy (anonymous):

5x1=5 5+1=6

OpenStudy (anonymous):

(x+1)(x+5) x+1=0 x+5=0 solve for x